### v1

# Product of Experts for Visual Generation

Abstract
Modern neural models capture rich priors and have complementary knowledge over shared data domains, e.g., images and videos. Integrating diverse knowledge from multiple sources—including visual generative models, visual language models, and sources with human-crafted knowledge such as graphics engines and physics simulators—remains under-explored. We propose a Product of Experts (PoE) framework that performs inference-time knowledge composition from heterogeneous models. This training-free approach samples from the product distribution across experts via Annealed Importance Sampling (AIS). Our framework shows practical benefits in image and video synthesis tasks, yielding better controllability than monolithic methods and additionally providing flexible user interfaces for specifying visual generation goals.  [^1] Project page: https://product-of-experts.github.io/.
1 Introduction
Visual generation requires diverse expertise. State-of-the-art image and video generation models have shown remarkable success in realistic appearance modeling, but still have limitations in producing physically accurate outputs or following complex user instructions. The former would benefit from domain expertise embedded in physics simulators, and the latter from semantic knowledge in large language models (LLMs) and visual language models (VLMs). However, training a single, monolithic model to absorb these varying knowledge sources (massive texts and visual datasets, simulation trajectories, etc.) is often impractical due to misaligned data distributions and training objectives.
We propose to address the challenge by integrating knowledge across diverse models at inference time. Specifically, we formulate visual generation in an inference-time Product of Experts (PoE) framework that aims to synthesize visual content leveraging specialized, pre-existing knowledge sources du2024compositional , including visual generative priors, discriminative rewards from LLMs/VLMs, and rule-based knowledge from physics simulators.
A naïve approach for incorporating multiple constraints is rejection sampling: generate candidate samples (or complete trajectories) using a base expert model, then discard failing ones. However, this is computationally prohibitive in high-dimensional visual synthesis due to vanishing acceptance rates. To improve sample efficiency, we utilize an annealed sampling procedure neal2001annealed with Sequential Monte Carlo (SMC) doucet2001introduction . Instead of inefficient post-hoc filtering, we propose to implement local (spatial for images and spatio-temporal for videos) updates from both generative and discriminative experts over many sampling steps, effectively steering the generation towards valid outputs.
The proposed framework can generate scenes better aligned with user inputs by explicitly integrating complementary expert knowledge sources without costly retraining. Its compositionally provides critical sample efficiency that enables image generation following complex text instructions. It also naturally yields a flexible user interface for defining complex generative goals by selecting and configuring experts, e.g., allowing users to specify the pose and motion trajectory of an object of interest, insert it into an existing image, and animate the full scene.
Overall, our contributions in this work are threefold:
  • i)
    We formulate visual synthesis tasks under constraints from heterogeneous generative and discriminative models in a unified Product-of-Experts (PoE) framework.
  • ii)
    We propose a practical, sample-efficient framework implementation with Annealed Importance Sampling (AIS) and Sequential Monte Carlo (SMC), with instantiated experts ranging across visual generative models, VLMs, and non-neural physics simulators.
  • iii)
    Our method enables flexible forms of user-instruction-following (texts, images, and low-level specifications such as object poses and trajectories) for image/video synthesis applications and achieves better user controllability and output fidelity compared to baseline methods.
2 Related Works
Compositional Generative Modeling.
Our work is related to recent work on compositional generative modeling, where multiple generative models are combined to jointly generate data samples (du2020compositional, ; garipov2023compositional, ; du2023reduce, ; mahajan2024compositional, ; du2024compositional, ; bradley2025mechanisms, ; thornton2025composition, ; gaudi2025coind, ). Compositional models have been previously applied in visual content generation (liu2022compositional, ; bar2023multidiffusion, ; zhang2023diffcollage, ; yang2024probabilistic, ; su2024compositional, ; li2022composing, ), but such applications typically focus on multiple homogeneous generator experts, or, in the case of li2022composing , a single generator expert with multiple discriminative experts. In contrast, our approach provides a general probabilistic framework through which many heterogeneous generative and discriminative experts can be jointly combined in generation.
Reward Steering from Discriminative Models.
Our work is further related to recent work steering generative models with discriminative models and reward functions grathwohl2019your ; dhariwal2021diffusion ; chung2022score ; song2023loss ; bansal2023universal ; he2023manifold ; ye2024tfg ; rout2025semantic . Methods typically use gradient descent on a discriminative reward grathwohl2019your ; dhariwal2021diffusion ; bansal2023universal , which assumes dense, differentiable gradients on the reward and are often approximate samplers that do not support sparse rewards models, e.g., VLMs. A recent branch of work has illustrated how Sequential Monte Carlo can be used for more exact reward steering (wu2023practical, ; zhao2024probabilistic, ). Our work builds on top of this formulation and generalizes it to sample from a composition of generative and discriminative experts.
Video Generation with Physical Simulators.
Our framework integrates both the data-driven knowledge from neural models and the domain-specific knowledge from non-neural models, e.g., physical simulators, allowing for more controllable, physically accurate video generation than end-to-end video generation models brooks2024video ; wan2025 , while alleviating the need for tedious 3D scene setups in traditional graphics rendering pipelines. Recent works have explored such direction typically convert physical simulator outputs into a specific type of conditional signals such as optical flow burgert2025go ; montanaro2024motioncraft ; yang2025towards and point tracking gu2025das , or rely on a single pre-trained video generation model liu2024physgen ; chen2025physgen3d ; tan2024physmotion for video-domain prior, while our framework allows for the integration of various signals available from physical simulation, providing improved flexibility and more complete specifications of controls.
3 The Product of Experts (PoE) Framework
We aim to generate complex scenes $x\in\mathbb{R}^{d}$ by combining the rich priors inside both generative models (such as autoregressive, diffusion models, and control signals converted from classical physics simulators) as well as discriminative models (such as VLMs), with examples detailed in Section 4. We formulate this composition of models in a probabilistic manner, where generative models are represented as probabilistic priors over scenes $\{p^{(i)}(\cdot)\}_{i=1}^{N}$. Discriminative models are represented as constraint functions, $\{r^{(j)}(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}\}_{j=1}^{M}$, each assigns a scalar score that represents the degree to which a scene $x$ matches the constraint encoded in the model. Each constraint function is converted into an unnormalized probability distributions through the Boltzmann distribution, $q^{(j)}:=e^{r^{(j)}}$.
To combine generative and discriminative models $\{p^{(i)}(\cdot)\}_{i=1}^{N}$ and $\{q^{(j)}(\cdot)\}_{j=1}^{M}$, we sample from the product distribution
$x\sim p(x):\propto\prod_{i=1}^{N}p^{(i)}(x)\prod_{j=1}^{M}q^{(j)}(x),$ (1)
where each generative expert $p^{(i)}(x)$ and discriminative expert $q^{(j)}(x)$ are defined over parts of the scene $x$. The product distribution $p(x)$ has high likelihood at scenes $x$ where each part of the scene is high likelihood under both $p^{(i)}(x)$ and $q^{(j)}(x)$ hinton2002training ; du2020compositional ; du2024compositional , allowing us to compose together the priors in both generative and discriminative models.
However, sampling from Eq. 1 is challenging, as we need to search for a sample $x$ that satisfies constraints across all experts. Below, we will discuss how we can efficiently and effectively sample from the Eq. 1 for high fidelity visual generation. We first discuss how we can sample from multiple generative experts in Section 3.1 and then discuss how to sample jointly from multiple discriminative and generative experts in Section 3.2, followed by a practical implementation in Section 3.3.
3.1 Sampling from Generative Experts
We first discuss how we can effectively sample from the product of a set of generative experts,
$x\sim p(x):\propto\prod_{i=1}^{N}p^{(i)}(x).$ (2)
To sample from the above distribution, we can use Markov chain Monte Carlo (MCMC) robert1999monte , where we can iteratively refine samples based on their likelihood under the product distribution. In discrete settings, a variant of Gibbs sampling may be used, while in continuous domains, Langevin sampling based on the gradient of log density of each distribution may be used welling2011bayesian .
A central issue, however, is that MCMC is a local refinement procedure, where it can take an exponentially long time for the sampling procedure to mix and find a high likelihood scene $x$ from the product distribution robert1999monte . To effectively sample from product of generative experts, we propose to use Annealed Importance Sampling (AIS) neal2001annealed , where we construct of a sequence of $T$ distributions, $\{p_{t}(x)\}_{t=1}^{T}$, where $p_{T}(x)$ corresponds to a smooth, easy-to-sample distribution and $p_{1}(x)$ corresponds to the desired product distribution defined in Eq. 2. To draw a sample from the product distribution $p_{1}(x)$, we first initialize a sample from $p_{T}(x)$ and iteratively run MCMC on the sample to sample from each intermediate distribution before we finally draw a sample from $p_{1}(x)$. Since intermediate distributions are smoother and easier to sample from, this enables us to more effectively find a high likelihood scene $x$ from the product distribution.
To construct intermediate probability distributions, in the continuous setting with diffusion (or flow) models, we can use the intermediate score functions learned by each composed diffusion model to define intermediate distributions. In particular, for each diffusion expert $p^{(i)}(x)$, the score functions $\{\nabla_{x}\log p_{t}^{(i)}(x)\}_{t=1}^{T}$ defines T separate distributions $p_{t}^{(i)}(x)$. We can then construct a sequence of annealed distributions $\{p_{t}(x)\}_{t=1}^{T}$, where
$p_{t}(x):\propto\prod_{i=1}^{N}p^{(i)}_{t}(x),$ (3)
and each distribution $p_{t}(x)$ is specified by the composed score function (Section 3.3), allowing us to use Langevin sampling to implement AIS of the product distribution.
While the main experiments on generative models in this paper use diffusion models, we can likewise define a sequence of intermediate distributions for other generative models. In the setting of autoregressive models, we can construct annealed distributions by considering the intermediate autoregressive marginals of the distribution $p(x_{1})$, $p(x_{1:2})$ and so forth, where $x_{1},x_{2}$ are data regions, or different tempered variants of the product distribution. We can then use Gibbs sampling to implement AIS of the product distribution.
Composing Conditional Generative Experts.
To improve the efficacy of sampling from the product distribution of generative experts, we can modify each generative expert to be conditionally dependent on the outputs of other experts. Specifically, we can represent the product distribution as
$x\sim p(x):\propto\prod_{i=1}^{N}p^{(i)}(x_{i}|x_{\text{pa}(i)}),$ (4)
where each $x_{i}$ refers to the part of a scene $x$ represented by the generative expert $p^{(i)}(\cdot)$, and $x_{\text{pa}(i)}$ refers to the other parts of a scene specified by other experts the expert is conditioned on.
Making each generative expert conditionally dependent on the values of other experts reduces the multi-modality in the distribution $p^{(i)}(\cdot)$, enabling more efficient MCMC sampling on the product distribution. We empirically illustrate the benefits of such a conditional generative expert formulation in Fig. 2, where sampling from the product of conditionally dependent generative models enables more effective sampling from the final product distribution.
3.2 Parallel Sampling with Discriminative Experts
Algorithm 1 Product of Expert (PoE) Sampling (Sketch)
1:Input: Annealing length $T$, initial distribution $p_{T}(\cdot)$, particle count $L$, MCMC step count $K$, generative experts $\{p_{t}^{(i)}(\cdot)\}_{i=1,t=1}^{N,T}$ with kernels $\mathcal{K}_{t\leftarrow t+1}(\cdot\mid\cdot)$ for MCMC initialization and $\mathcal{K}_{t}(\cdot\mid\cdot)$ for MCMC steps, discriminative experts with log probabilities $\{r^{(j)}(\cdot)\}_{j=1}^{M}$.
2:Initialize: Sample $x^{(l)}\sim p_{T}(\cdot)$, $\forall l=1,\cdots,L$.
3:for $t=T-1$ to $1$ $\triangleright$ Transition to the next annealed distribution
4:     for $l=1$ to $L$ do $\triangleright$ Parallel sampling
5:         $x^{(l)}\leftarrow\mathcal{K}_{t\leftarrow t+1}(\cdot\mid x^{(l)})$triangleright MCMC initialization
6:         for $k=1$ to $K$ do
7:              $x^{(l)}\leftarrow\mathcal{K}_{t}(\cdot\mid x^{(l)})$triangleright MCMC step
8:         end for
9:     end for
10:     Resample samples with log weights $\sum_{j=1}^{M}r_{t}^{(j)}(x^{(l)})$ where $r_{t}^{(j)}(x^{(l)})\approx r^{(j)}(\hat{x}^{(l)})$.
11:end for
12:$l^{*}\leftarrow\max_{l}\sum_{j=1}^{M}r^{(j)}(x^{(l)})$
13:Output: $x^{(l^{*})}$
Next, we discuss how we can modify the annealed sampling procedure in Section 3.1 to sample from the full product distribution, including discriminative experts
$x\sim p(x):\propto\prod_{i=1}^{N}p^{(i)}(x)\prod_{j=1}^{M}q^{(j)}(x).$ (5)
Similar to the previous section, we can define a sequence of intermediate sampling distributions
$p_{t}(x):\propto\prod_{i=1}^{N}p^{(i)}_{t}(x)\prod_{j=1}^{M}q^{(j)}_{t}(x).$ (6)
To implement AIS with these intermediate distributions, one simple approach is to use importance sampling. Similar to the previous section, we can use MCMC sampling to sample from the product of the intermediate generative expert distributions from Eq. 3. After generating a total of $L$ samples from this generative product distribution, we draw a weighted random sample from these $L$ samples, with each sample has an importance weight $\prod_{j=1}^{M}q^{(j)}_{t}(x)$, corresponding to the likelihood of the scene under the product of the discriminative experts.
An issue with this simple procedure is that each of the $L$ samples drawn from the MCMC may be correlated with each other, since MCMC is slow at mixing and may not cover the entire generative product distribution. To improve the coverage of the $L$ samples drawn from MCMC, we use a parallel SMC sampler doucet2001introduction , where we maintain $L$ particles over the course of annealed sampling. At each intermediate annealed distribution, we run MCMC on each particle to draw samples from the product of the generative distributions, and then weigh and resample particles based on their likelihood under the product of the discriminative experts.
Our overall algorithm for sampling from Eq. 5 requires only black-box access to each discriminator, requiring only knowledge of the likelihood a discriminator assigns to a sample $x$ and not additional information such as the gradient of the log-likelihood of the discriminator. For settings where we do have more detailed information about the form of a discriminative expert, we can then directly treat the discriminative expert as a generative expert, as in Section 3.1, and use approaches such as gradient-based MCMC sampling to efficiently sample from the composed expert distribution.
3.3 Framework Instantiation
Figure 1: Application on Image Object Insertion. Given a graphics engine configurations (top row), a textual description, and an input image (first column), our method produces visually plausible edits with objects inserted.
We now introduce a specific implementation of the framework as summarized in Algorithm 1, which describes the annealed distribution and expert instantiations primarily used in the paper.
Constructing Annealed Distribution with Flow Models.
We construct our annealed generative expert distributions $p_{t}(x)$ (Eq. 3) using flow models. Each $p_{t}^{(i)}(x)$ is generated with a predicted velocity $v^{(i)}_{t}(x)$, which can be converted to score $\nabla_{x}\log p_{t}^{(i)}(x)$ through linear transformations. To propagate a sample from one annealed distribution $p_{t+1}(x)$ to the next $p_{t}(x)$, we use one Euler step in the flow ODE using the composed velocity $v_{t}(x)=\sum_{i}v^{(i)}_{t}(x)$ of the generative expert distributions, corresponding to a transition kernel $\mathcal{K}_{t\leftarrow t+1}(x^{\prime}\mid x):=\delta(x^{\prime}-(x+v_{t}(x)\Delta t))$ where $\delta$ denotes the Dirac delta function and $\Delta t$ is the discretization step size. While this integration procedure does not draw a sample from $p_{t}(x)$, it serves as an effective initialization for proper MCMC sampling from $p_{t}(x)$. We then use Langevin dynamics on the composed score $s_{t}(x)=\sum_{i}\nabla_{x}\log p_{t}^{(i)}(x)$ to sample from $p_{t}(x)$, i.e., $\mathcal{K}_{t}(x^{\prime}\mid x):=\mathcal{N}(x^{\prime};x+\frac{\kappa^{2}}{2}s_{t}(x),\kappa^{2}I)$ with step size $\kappa$. Constructions with autoregressive models are deferred to Appendix C.
Conditional Sampling.
We implement conditional generative experts $p_{t}^{(i)}(x_{i}\mid x_{\text{pa}(i)})$ by modifying the original generative experts $p_{t}^{(i)}(x_{i})$. In particular, we modify the generative flow $v^{(i)}_{t}(x_{i})$ to take into account the predicted flow at parent regions
$v_{t}^{(i)}(x_{i}\mid x_{i^{\prime}})\approx v^{(i)}_{t}(x_{i})-w\sum_{i^{\prime}\in\text{pa}(i)}\nabla_{x_{i}}\|v^{(i)}_{t}(x_{i})-\text{stopgrad}(v^{(i^{\prime})}_{t}(x_{i^{\prime}}))\|_{2}^{2},$ (7)
where $w$ is the learning rate. Justifications of this approximation are deferred to Section D.2.
Discriminative Experts.
To define each annealed discriminative expert $q^{(j)}_{t}(x)$ with reward function $r^{(j)}(x)$, we define its intermediate counterpart $r_{t}^{(j)}(x):=r^{(j)}(\hat{x})$, where $\hat{x}$ is the endpoint of $x$ predicted using the flow model $v_{t}(x)$. We use pre-trained VLMs as reward functions in this work.
Figure 2: Baseline Comparisons on Image Object Insertion. Our method better adheres to input object pose configurations compared to RF-Solver. The 4-th column shows conditional sampling improves visual fidelity.
Methods Controllability Image Quality Text Alignment
MSE (bg) ($\downarrow$) LPIPS (fg) ($\downarrow$) GPT-4o ($\uparrow$) Aesthetic ($\uparrow$) GPT-4o ($\uparrow$) ImageReward ($\uparrow$) GPT-4o ($\uparrow$)
RF-Solver $1.619$±$\pm 0.605$ $\underline{0.178}$±$\pm 0.061$ $0.518$±$\pm 0.168$ $\mathbf{0.618}$±$\pm 0.070$ $0.662$±$\pm 0.103$ $0.948$±$\pm 0.777$ $0.438$±$\pm 0.179$
Ours $\mathbf{1.429}$±$\pm 0.602$ $\mathbf{0.065}$±$\pm 0.019$ $\mathbf{0.827}$±$\pm 0.073$ $0.596$±$\pm 0.061$ $\mathbf{0.783}$±$\pm 0.082$ $\mathbf{1.175}$±$\pm 0.666$ $\mathbf{0.817}$±$\pm 0.129$
No Cond $\underline{1.511}$±$\pm 0.661$ $\mathbf{0.065}$±$\pm 0.018$ $\underline{0.727}$±$\pm 0.141$ $\underline{0.599}$±$\pm 0.059$ $\underline{0.718}$±$\pm 0.120$ $\underline{1.142}$±$\pm 0.668$ $\underline{0.675}$±$\pm 0.180$
Table 1: Image Object Insertion. Our method achieves better foreground, background, and textual alignment and comparable visual quality compared to RF-Solver; $\pm$ denotes standard deviation.
4 Experiments
We now show results applying the PoE framework to image/video synthesis tasks. Expert configurations can adapt to the desired, task-dependent input granularity, including high-level text instructions (Sections 4.2, 4.3 and 4.4) and precise, low-level pose controls (Sections 4.2 and 4.3), and leverage expert knowledge ranging across natural image and video priors (Sections 4.2, 4.3 and 4.4), precise physics rules from physical simulators (Section 4.3), and semantic visual understandings from VLMs (Section 4.4).
4.1 Implementation Details
As described in Eq. 4, a generative expert can be defined over the entire scene $x\in\mathbb{R}^{d}$ or a region of the scene $x_{i}\in\mathbb{R}^{d^{(i)}}$ with $d^{(i)}\leq d$, i.e., the expert has prior over only part of the scene, e.g., a local spatial region in Section 4.4. In these cases, we zero-pad the velocity prediction from the generative expert from $\mathbb{R}^{d^{(i)}}$ to $\mathbb{R}^{d}$ before velocity composition (Algorithm 1, line 5).
In other cases, one may restrict the influence of a generative expert to update only a scene region, even when the expert is defined over the full scene $x\in\mathbb{R}^{d}$, e.g., when the expert only has knowledge over foreground object informed by the graphics engine or physics simulators in Sections 4.2 and 4.3. We apply a local (spatial for images and spatio-temporal for videos) weight $\lambda^{(i)}(x)\in\left[0,1\right]^{d}$ on the velocity prediction $v_{t}^{(i)}(x)$ in these cases, where $\lambda^{(i)}(x)$ is the Gaussian-blurred foreground mask.
4.2 Graphics-Engine-Instructed Image Editing
Task Specifications.
We apply our framework on the task of image editing where user provides an image to be edited, a 3D asset posed in a graphics engine, Genesis [^2] https://github.com/Genesis-Embodied-AI/Genesis, to precisely describe where to insert the asset, and a text description describing higher-level information such as object materials (“metal”) and semantics (“coke can”). We deploy two generative experts: a depth-to-image model, FLUX.1 Depth [dev] flux2024 , and an inpainting model, FLUX.1 Fill [dev] flux2024 . The former provides necessary constraints on the outputs to ensure objects inserted follow input pose specifications, and the latter provides a natural image prior to produce realistic visual outputs.
Evaluation.
The task is evaluated on a dataset with 10 natural images paired with 3 object inputs, resulting in 30 scenes in total, with examples shown in Fig. 1. Evaluation metrics cover three major visual aspects, including controllability, image quality, and text alignment. Specifically, we evaluate the MSE distance between generation outputs and input images on background pixels to evaluate background preservation, and the LPIPS zhang2018perceptual distance between outputs and graphics engine’s RGB renderings on foreground pixels to measure the fidelity to input objects. We report aesthetic score using the LAION-Aesthetic predictor [^3] https://github.com/LAION-AI/aesthetic-predictor and measure text alignment with ImageReward xu2023imagereward . Finally, for each evaluation aspect, we query GPT-4o for automatic evaluation. See details in Section F.1.
Results.
We compare our method with an inversion-based image editing baseline, RF-Solver wang2024rfsolver , which receives the same input image and a target text prompt, e.g., “a metal box standing on the ground”, which is inherently coarse and therefore does not produce results conforming to the graphics engine inputs, and it sometimes fail to faithfully preserves input background as shown in Fig. 2. Table 1 suggests that our method achieves better controllability on the task.
Ablation on Conditional Sampling.
Section 3.1 introduced the conditionally sampling strategy, which is crucial for efficient sampling from the product distribution. Removing the updates in Eq. 7 (no-cond in Table 1) decreases performance and results in less harmonized visual outputs (Fig. 2).
Figure 3: Application on Physical-Simulation-Instructed Video Generation. Given an input image and an input sequence from physical simulators describing precise object motions, our method generates videos aligned with input object motions while synthesizing natural video content for non-foreground regions.
Figure 4: Baseline Comparisons on Physics-Simulator-Instructed Video Generation. All frames are processed in grayscale and overlaid with tracking trajectories predicted by SpaTracker xiao2024spatialtracker for visualization.
Figure 5: Baseline Comparisons on PhysGen3D Data. Our method improves the visual quality compared to the input simulation and adheres more closely to the input object motions compared to baselines.
Methods Controllability Video Quality Semantic Alignment
LPIPS ($\downarrow$) GPT-4o ($\uparrow$) Smooth ($\uparrow$) Aesthetic ($\uparrow$) Imaging ($\uparrow$) GPT-4o ($\uparrow$) ViCLIP ($\uparrow$) GPT-4o ($\uparrow$)
Object-Centric Simulation Input
Traj2V $0.109$±$\pm 0.033$ $0.638$±$\pm 0.172$ $\underline{0.993}$±$\pm 0.003$ $0.534$±$\pm 0.071$ $0.570$±$\pm 0.154$ $0.725$±$\pm 0.179$ $0.258$±$\pm 0.041$ $0.757$±$\pm 0.141$
Depth2V $\mathbf{0.103}$±$\pm 0.031$ $\underline{0.650}$±$\pm 0.144$ $0.984$±$\pm 0.015$ $0.534$±$\pm 0.069$ $0.581$±$\pm 0.155$ $0.750$±$\pm 0.126$ $\underline{0.261}$±$\pm 0.041$ $0.775$±$\pm 0.130$
Image2V $0.111$±$\pm 0.021$ $\mathbf{0.708}$±$\pm 0.104$ $0.978$±$\pm 0.025$ $\mathbf{0.553}$±$\pm 0.057$ $\underline{0.611}$±$\pm 0.101$ $\underline{0.775}$±$\pm 0.099$ $0.255$±$\pm 0.036$ $\underline{0.788}$±$\pm 0.079$
Ours $\underline{0.104}$±$\pm 0.030$ $\mathbf{0.708}$±$\pm 0.104$ $\mathbf{0.994}$±$\pm 0.003$ $\underline{0.549}$±$\pm 0.068$ $\mathbf{0.625}$±$\pm 0.117$ $\mathbf{0.842}$±$\pm 0.132$ $\mathbf{0.270}$±$\pm 0.038$ $\mathbf{0.817}$±$\pm 0.080$
Full-Scene Simulation Input (PhysGen3D dataset)
PhysGen3D $0.438$±$\pm 0.099$ $\mathbf{0.995}$±$\pm 0.002$ $0.571$±$\pm 0.114$ $0.673$±$\pm 0.046$ $0.513$±$\pm 0.117$ $0.224$±$\pm 0.026$ $0.588$±$\pm 0.117$
Inversion $0.237$±$\pm 0.084$ $0.263$±$\pm 0.122$ $\underline{0.993}$±$\pm 0.002$ $0.468$±$\pm 0.097$ $0.410$±$\pm 0.144$ $0.250$±$\pm 0.206$ $0.202$±$\pm 0.058$ $0.400$±$\pm 0.206$
Depth2V $\underline{0.164}$±$\pm 0.065$ $\underline{0.550}$±$\pm 0.173$ $\underline{0.993}$±$\pm 0.003$ $\underline{0.579}$±$\pm 0.067$ $\underline{0.680}$±$\pm 0.081$ $\underline{0.662}$±$\pm 0.216$ $\mathbf{0.242}$±$\pm 0.030$ $\underline{0.788}$±$\pm 0.108$
Image2V $0.242$±$\pm 0.077$ $0.525$±$\pm 0.192$ $0.986$±$\pm 0.010$ $\mathbf{0.594}$±$\pm 0.074$ $0.662$±$\pm 0.052$ $0.638$±$\pm 0.245$ $0.236$±$\pm 0.028$ $\underline{0.788}$±$\pm 0.105$
Ours $\mathbf{0.136}$±$\pm 0.047$ $\mathbf{0.587}$±$\pm 0.220$ $\underline{0.993}$±$\pm 0.003$ $0.575$±$\pm 0.071$ $\mathbf{0.688}$±$\pm 0.045$ $\mathbf{0.763}$±$\pm 0.132$ $\underline{0.239}$±$\pm 0.038$ $\mathbf{0.825}$±$\pm 0.141$
Table 2: Physics-Simulator-Instructed Video Generation.
4.3 Physical-Simulator-Instructed Video Generation
Task Specifications.
The setup from Section 4.2 can be directly extended to dynamic scenes for video generation, with object poses and dynamics (e.g., a ball bouncing, Fig. 3 top row) specified by a physical simulator, an input image, and a textual scene description. We use Genesis as the simulator, and use two generative experts: an image-to-video model, Wan 2.1 wan2025 , and a depth-to-video model [^4] We use the released checkpoint on Wan 2.1 from https://github.com/aigc-apps/VideoX-Fun..
Evaluation.
We evaluate our method and several monolithic models, including image-to-video (Image2V), depth-to-video (Depth2V), and trajectory-to-video (Traj2V), all using the same Wan2.1 (14B) model as backbone, on a dataset consisting of 3 simulation inputs paired with 4 input images each, giving a total of 12 scenes with examples in Fig. 3. To obtain the initial frame for all methods, we edit the input image with our method described in Section 4.2 to insert the simulated object. Depth maps required as method inputs are rendered from the physical simulator, and input trajectory maps are computed using object centroids. We report metrics similar to Section 4.2, with additional ones adopted from VBench huang2024vbench : motion smoothness scores introduced in VBench huang2024vbench , MUSIQ ke2021musiq for imaging quality, and ViCLIP wang2023internvid for overall semantic consistency with input text prompts.
In the above setting, simulation inputs only describe foreground objects. We further evaluate on cases where full-scene simulation is available. We use the simulation results from 9 released scenes from PhysGen3D chen2025physgen3d , a recently proposed method that reconstructs and animates images in physical simulators, as evaluation inputs. The full-scene simulation is closely aligned with natural videos, and therefore we use the noise vector from flow ODE inversion for particle initialization as opposed to random noise (Algorithm 1, line 2). We further compare with flow inversion baseline for this setting.
Results.
When only object-centric simulation is available (Table 2, top half), depth-to-video and trajectory-to-video alone fail to capture rich non-foreground-object motions, such as cars moving and birds running in Fig. 4. In particular, they tend to compensate for the object’s downward motion with upward camera motions, as visualized by the tracking trajectories in Fig. 4. With full-scene simulation as inputs (Table 2, bottom half), our method improves the video quality over the input (e.g., the apple’s black regions in Fig. 5) while remaining faithful. On the other hand, the depth-conditioned baseline deviates from the input simulation as shown in Fig. 5. In both cases, the image-to-video model does not adhere to the input simulation due to the coarseness of text descriptions.
Figure 6: Text-to-Image Generation using regional generative experts and VLM as discriminative expert.
4.4 Text-to-Image Generation with Regional Composition
We further demonstrate the efficacy of distribution composition on the task of text-to-image generation, where we first query an LLM to parse an input text prompt into spatial bounding boxes with regional prompts (Fig. 6 left), aiming to compose the distributions from text-to-image model (FLUX.1 [dev]) conditioned on local text prompts, one for each local region, as generative experts, and from VQAScore lin2024evaluating as a discriminative expert, which gives a scalar value predicting the probability that its input image is aligned with the input text prompt. We compare our methods using $L=16,4$ particles and $1$ particle (which amounts to not using the discriminative expert), a variant using no generative composition (equivalent to best-of-N sampling for vanilla FLUX) while matching the compute budget for 16 particles, and with the vanilla FLUX backbone. In the examples shown in Fig. 6, our compositional approach achieves better text alignment compared to the baselines, and the discriminative expert provides crucial knowledge, e.g., counting, to steer the output distribution.
5 Conclusion
We have introduced a Product of Experts (PoE) framework for composing knowledge from heterogeneous sources at inference time for visual generation, and proposed an efficient sampling algorithm leveraging Sequential Monte Carlo. Our training-free approach effectively integrates visual generative priors, VLM guidance, and physics-based constraints. We demonstrated that this method generates images and videos with higher complexity and improved controllability compared to monolithic approaches, providing a promising direction for the next generation of visual generation models.
Acknowledgments.
This work was in part supported by ONR N00014-23-1-2355, ONR YIP N00014-24-1-2117, ONR MURI N00014-22-1-2740, and NSF RI #2211258. YZ is in part supported by the Stanford Interdisciplinary Graduate Fellowship.
Appendix A Overview
The supplementary material contains discussions (Appendix B), extended descriptions of the framework (Appendices C and D), and additional experimental results (Appendix E) and details (Appendix F).
Appendix B Discussions
Compute Requirement.
All image experiments are run using 1 NVIDIA H200 GPUs and all video experiments with 2 NVIDIA H200 GPUs. Generating one sample in Sections 4.2, 4.3, 4.4 and C takes approximately 4min, 5min, 4min, and 30min, respectively.
Limitations.
The proposed method requires extra computation due to intermediate MCMC steps and parallel sampling compared to vanilla feedforward inference of the image or video generative model backbones. Furthermore, the sampling efficacy relies on the assumption that expert predictions are reasonably compatible, and strong incompatibilities might require other sampling techniques or training. This work proposes a framework and provides empirical validation with small number of experts and we leave large-scale PoE sampling as future work.
Societal Impact.
We believe outputs of the proposed framework does not directly possess negative societal impact. The underlying technology enable more content creation tools and benefits understanding model behaviors. However, highly controllable generation could be exploited to create misinformation. While our work does not inherently use sensitive data, we are aware of such risks and closely follow ethical guidelines in the community to help mitigate these risks.
Appendix C Framework Instantiation with Autoregressive Experts
As described in Section 3.3, instantiating the framework requires constructing the annealed distributions $\{p_{t}(x)\}_{t=1}^{T}$in Eq. 3, and defining the corresponding transition kernel from $p_{t}(x)$ to $p_{t-1}(x)$. The main paper uses flow ODE discretization to construct the annealed distributions.
An alternative construction is to use autoregressive marginals $p_{t}^{(i)}(x):=\bar{p}^{(i)}(x_{1:t})$ such that $x_{1:T}=x$. Note that we flip the order of $t$ to traverse from $1$ to $T$ for notation convenience. We use bar notations for pre-trained models in the following discussions. Given an autoregressive model that allows for sampling from $\bar{p}(x_{1:t+1}\mid x_{1:t},\theta_{i})$, Algorithm 1 can be implemented with $\mathcal{K}_{t\leftarrow t-1}(x^{\prime}\mid x):=\prod_{i=1}^{N}\bar{p}^{(i)}(x_{1:t}^{\prime}\mid x_{1:t-1})$ and $\mathcal{K}_{t}$ as Gibbs sampling.
The following are empirical results under this construction. We use two generative experts: a frame-wise autoregressive video generation model, FramePack zhang2025framepack , and a graphics engine, Genesis, which defines a Gaussian distribution centered on a video sequence of the physics simulator’s RGB rendering overlayed with the input initial frame, visualized as “Input Sim” in Fig. 7. In this case, $x_{1:t}$ represents the first $t$ frames of a video $x$, and the product distribution is defined as $p(x)\propto e^{-w\|x-c^{\text{simulation}}\|^{2}}p^{\text{image2video}}(x\mid c^{\text{text}},c^{\text{image}})$ with some constant $w\in\mathbb{R}$. We approximate sampling from this distribution with first sampling from $p^{\text{image2video}}$, then do 8 gradient updates with respect to $\mathcal{L}_{2}$ loss $\|x-c^{\text{simulation}}\|^{2}$.
Qualitative results for the application from Section 4.3 and comparisons with vanilla image-to-video inference using the same backbone model, HunyuanVideo kong2024hunyuanvideo  [^5] Pre-trained weights from https://huggingface.co/lllyasviel/FramePack_F1_I2V_HY_20250503., are shown in Fig. 7, with $81$ frames per sequence and spatial resolution $512\times 768$. We set $L=1$, with 1 MCMC sampling step per iteration for this experiment. Results from our method follow the input physics simulation while synthesizing motions from other subjects in the scenes following the natural video prior from the backbone model.
Figure 7: Framework Instantiation with Autoregressive Annealed Distributions. Given an input first frame (first column), a text prompt (at the bottom of each subfigure), and the RGB rendering from a physics simulator overlayed with the input first frame (top row of each subfigure), our method generates natural videos following the object motion specified by physics simulations. In comparison, the baseline image-to-video model does not produce the desired object motion.
Appendix D Framework Details
D.1 Flow Models
Consider the $i$-th generative expert, which is a flow model that defines a time-dependent continuous probability path $\bar{p}^{(i)}(x;\bar{t}),x\in\mathbb{R}^{d},\bar{t}\in\left[0,1\right]$, with boundary conditions $\bar{p}^{(i)}(x;0)=\mathcal{N}(0,I)$ and $\bar{p}^{(i)}(x;1)\approx p_{\text{data}}^{(i)}(x)$. Each flow model is associated with a scheduler $\bar{\alpha},\bar{\sigma}:\left[0,1\right]\rightarrow\mathbb{R}$ that defines a conditional probability path $\bar{p}^{(i)}(x\mid y;\bar{t})=\mathcal{N}(x\mid\bar{\alpha}(\bar{t})y,\bar{\sigma}(\bar{t})I),y\sim p_{\text{data}}^{(i)}(y)$, and is trained to predict a velocity field $\bar{v}^{(i)}(x;\bar{t},\theta_{i})$ that generates the marginalized path $\bar{p}^{(i)}(x;\bar{t})$, where $\theta_{i}$ denotes model parameters. While the training-time scheduler $(\bar{\alpha},\bar{\sigma})$ can differ across experts, they can be aligned during inference time via a scale-time transformation lipman2024flow .
Given a time remapping function $\xi:\{T,T-1,\cdots,1\}\rightarrow\left[0,1\right]$, one can construct the per-expert product component from Eq. 3 via discretizing the path $\bar{p}^{(i)}(x;\bar{t})$ into
$p^{(i)}_{t}(x):=\bar{p}^{(i)}(x;\xi(t)).$ (8)
We require that $\xi$ is monotonic and $\xi(T)=0,\xi(1)=1$. By construction, $p^{(i)}_{T}(x)=\mathcal{N}(0,I)$ is easy to sample from, and $p^{(i)}_{1}(x)=p_{\text{data}}^{(i)}(x)$ encapsulates the prior data distribution of the expert model, fulling our motivation in Section 3.1. For all experiments, we define $\xi$ to be the Euler discretization esser2024scaling ; euler1792institutiones , with additional alignment to the DDPM scheduler ho2020denoising for experiments in Section 4.4.
Line 5 from Algorithm 1 is computed as follows:
$v_{t}^{(i)}(x)$ $=\bar{v}^{(i)}(x;\xi(t),\theta_{i}),$
$\alpha_{t}$ $=\bar{\alpha}(\xi(t)),$
$\sigma_{t}$ $=\bar{\sigma}(\xi(t)),$
$s_{t}^{(i)}(x)$ $=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x-\frac{\dot{\sigma_{t}}\sigma_{t}\alpha_{t}-\dot{\alpha}_{t}\sigma^{2}_{t}}{\alpha_{t}}v_{t}^{(i)}(x),$
$v_{t}(x)$ $=\sum_{i}\lambda^{(i)}(x)v^{(i)}(x),$
$s_{t}(x)$ $=\sum_{i}\lambda^{(i)}(x)s^{(i)}(x),$
$\mathcal{K}_{t\leftarrow t+1}(x^{\prime}\mid x)$ $=\delta(x^{\prime}-(x+v_{t}(x)(\xi(t-1)-\xi(t)))),$
$\mathcal{K}_{t}(x^{\prime}\mid x)$ $=\mathcal{N}(x^{\prime};x+\frac{\kappa^{2}}{2}s_{t}(x),\kappa^{2}I),$
where $\lambda^{(i)}(x)$ is defined in Section 4.1, with all entries of $\sum_{i}\lambda^{(i)}(x)$ to be $1$.
D.2 Approximation for Conditional Sampling
We explain Eq. 7 below. Let $x_{i}$ be any expert, we then approximate the distribution $p(x_{i}\mid x_{\text{pa}(i)})$ as
$p(x_{i}\mid x_{\text{pa}(i)})\propto p(x_{i})p(x_{\text{pa}(i)}\mid x_{i})=p(x_{i})\prod_{i^{\prime}\in\text{pa}(i)}p(x_{i^{\prime}}\mid x_{i}),$ (9)
where parent regions $x_{i^{\prime}}$ are assumed to be conditionally independent given $x_{i}$. To model each conditional distribution $p(x_{i^{\prime}}\mid x_{i})$, we define a Gaussian distribution of the deviation of the flow vector predicted at $x_{i^{\prime}}$ and $x_{i}$, where $p(x_{i^{\prime}}\mid x_{i})\propto e^{-w\|v(x_{i^{\prime}})-v(x_{i})\|^{2}}$ with a constant $w\in\mathbb{R}$. Under this parametric distribution, when $x_{i^{\prime}}$ and $x_{i}$ have consistent flow predictions, the likelihood is high (which is reasonable as this indicates both $x_{i^{\prime}}$ and $x_{i}$ are mutually compatible). We can convert the probability expression in Eq. 9 to a corresponding score function, resulting in Eq. 7.
In this work, $\text{pa}(i)$ is defined as the generative expert defining a natural data distribution for global $x\in\mathbb{R}^{d}$, e.g., a text-conditioned generative model for the full image/video $x$. Each generative expert is additionally conditioned on a context signal $c_{i}\in\mathbb{R}^{d_{i}^{\text{context}}}$, e.g., texts or depth maps. Note that the framework allows for different types of conditions across experts, enabling flexible control handles that are typically application-dependent.
Appendix E Extended Experimental Results
E.1 Qualitative Video Results
Please view the “index.html” in the supplementary material folder using a web browser.
E.2 Intermediate Sample Visualizations
Figure 8: Visualizations of Parallel Sampling with Discriminative Experts. Discriminative expert (VQAScore) scores are annotated on the top-left corners; score underlining means the sample proceeds to the next annealed distributions (with smaller $t$). Images with the same boundary color share the same initial seed.
The parallel sampling process is visualized in Fig. 8. For each scene from Section 4.4, we use an LLM to convert a text prompt into a layout of bounding boxes following the procedure described in Section F.4. The input text prompts and resulting layouts and regional prompts are visualized in Fig. 6. Then, we follow Algorithm 1 and initialize $L=4$ particles and annealed schedule length $T=28$, which are reweighed at three key timesteps $t=21,19,1$ using the discriminative expert, VQAScore lin2024evaluating , which measures image-text alignment with a probability score. At $t=21,19$, we binarize the scores with the threshold set to be the median score of the current 4 particles, discard the two below the threshold, and duplicate the remaining two high-score particles two times to maintain 4 particles in each round. At $t=1$, we keep the highest-scoring particle as the final output sample.
Fig. 8 shows end-point prediction, $\hat{x}^{(l)}=\frac{\sigma_{t}}{\dot{\alpha}_{t}\sigma_{t}-\dot{\sigma}_{t}\alpha_{t}}v_{t}(x^{(l)})-\frac{\dot{\sigma}_{t}}{\dot{\alpha}_{t}\sigma_{t}-\dot{\sigma}_{t}\alpha_{t}}x^{(l)}$, to visualize the intermediate samples from the annealing process from $t=28$ to $t=1$. VQAScore measurements (higher means more aligned) are overlayed on the top-left corner for each sample. In the proposed framework, integrating multiple generative experts allows for composing local, simpler regional prompts, construction of the annealed distributions allows for iterative refinement of samples, and parallel sampling further integrates the knowledge from the discriminative expert and improves output accuracy, as evidenced in the comparisons from Fig. 6.
Appendix F Implementation Details
F.1 Automatic Evaluation with VLMs
We follow the protocol proposed in PhysGen3D chen2025physgen3d to run automatic evaluation for Sections 4.4 and 4.3 using GPT-4o hurst2024gpt4o . For each scene, all methods’ output images, or uniformly sampled 10 frames when outputs are videos, are sent to GPT-4o together with template prompts as specified in PhysGen3D chen2025physgen3d . GPT-4o is asked to give scores from 00 to $1$ for each video on all metrics. The scores for the same metric are averaged across all scenes.
F.2 Graphics-Engine-Instructed Image Editing
In Section 4.2, the product distribution is defined as $p(x)\propto p^{\text{depth2image}}(x\mid c^{\text{text}},c^{\text{depth}})p^{\text{imagefill}}(x\mid c^{\text{text}},c^{\text{image}},c^{\text{mask}})$. We use $T=28,L=1$, with 1 MCMC sampling step and 2 conditional sampling updates (Eq. 7) per iteration, and spatial resolution $960\times 1664$.
F.3 Physical-Simulator-Instructed Video Generation
In Section 4.3, the product distribution is defined as $p(x)\propto p^{\text{depth2video}}(x\mid c^{\text{text}},c^{\text{depth}})p^{\text{image2video}}(x\mid c^{\text{text}},c^{\text{image}})$. We use $T=30,L=1$ with 1 MCMC sampling step per iteration. We use no conditional sampling updates for this task due to GPU memory constraint. Each generated video has $81$ frames and spatial resolution $480\times 832$.
F.4 Text-to-Image Generation with Regional Composition
Text-to-Layout.
We use an LLM (GPT-o1-mini) to convert an input text prompt into a global caption, a list of region captions, and a layout specified by bounding boxes through the following steps.
We query LLM to separate the input text prompt into a list of main regions, each associated with a region caption. Then we pass the original prompt along with the separated list of region captions and ask it to assign a spatial box to each region, specified as a tuple of percentages: (start_x, end_x, start_y, end_y), where each value is in the range $[0,100]$. We instruct that these boxes are intended to form a plausible image layout that must satisfy all spatial constraints described in the prompt. We instruct to make each box as large as possible, subject to the following constraints: the width and height must each be at least 40; the width must be between 0.6×0.6$\times$0.6 × and 1.5×1.5$\times$1.5 × the height; neighboring boxes should overlap unless the prompt explicitly specifies that objects should be separated; however, the overlap between any two boxes must not exceed $50\%$.
Layout-to-Image Generation.
We use regional generative experts and discriminative experts for layout-to-image generation, or $p(x)\propto\left(\prod_{i}p^{\text{text2image}}(x_{i}\mid c^{\text{text}})\right)q^{\text{VLM}}(x\mid c^{\text{text}})$, as described in Section 4.4. Prompts from Figs. 8 and 6 are taken from GenAI-Bench li2024genai . In Section 4.4 and Fig. 8, we use $T=28$ and image resolution $1024\times 1024$, with $1$ MCMC sampling step per iteration.